# Toward Robust Hyper-Detailed Image Captioning: A Multiagent Approach and Dual Evaluation Metrics for Factuality and Coverage

### v2

@@ Abstract
Multimodal large language models (MLLMs) excel at generating highly detailed captions but often produce hallucinations. Our analysis reveals that existing hallucination detection methods struggle with detailed captions. We attribute this to the increasing reliance of MLLMs on their generated text, rather than the input image, as the sequence length grows. To address this issue, we propose a multiagent approach that leverages LLM-MLLM collaboration to correct given captions. Additionally, we introduce an evaluation framework and a benchmark dataset to facilitate the systematic analysis of detailed captions. Our experiments demonstrate that our proposed evaluation method better aligns with human judgments of factuality than existing metrics and that existing approaches to improve the MLLM factuality may fall short in hyper-detailed image captioning tasks. In contrast, our proposed method significantly enhances the factual accuracy of captions, even improving those generated by GPT-4V. Finally, we highlight a limitation of VQA-centric benchmarking by demonstrating that an MLLM’s performance on VQA benchmarks may not correlate with its ability to generate detailed image captions.
@@ 3.3 Evaluation Methods
To validate this evaluation method, we use the DOCCI dataset (Onoe et al., 2024), which contains human-annotated detailed image captions. Specifically, for each image in a subset of the dataset, we prepare the following four types of captions (details provided in Appendix F):
@@ 4.1 Experimental Setup
We employ our proposed factuality and coverage metrics, along with CLAIR, all based on GPT-4o, to evaluate the generated captions. To ensure robust evaluation, we summarize the captions (Ge et al., 2024) generated from five different prompts using LLaMA-3-8B. The only CapMAS hyperparameter, , is tuned on a validation set of five examples sampled from the DCI dataset (Urbanek et al., 2024). The prompt templates are provided in Appendix F.
@@ 4.4 Comparison with Other Methods
Various methods have been proposed to mitigate hallucinations in MLLMs, and they have primarily been validated on VQA and simple captioning benchmarks. We compare CapMAS with two recent decoding-based methods (VCD and OPERA), two corrector-based methods (LURE and Volcano), and one training-based method (LRV) from the perspective of detailed image captioning. All methods, except for LRV and LURE, use LLaVA-v1.5-7B, while the LRV and LURE methods employ the MiniGPT-4 model (Zhu et al., 2023) as provided by their respective authors. For reference, VisualFactChecker (VFC) (Ge et al., 2024) is also a pipeline composed of pre-trained models that revise initial captions, similar to our approach. However, the inability to reproduce VFC, as its authors have not provided the necessary resources for reproduction, prevents a direct comparison with our method. Nonetheless, we can infer that our method outperforms VFC in terms of both applicability and performance because 1) VFC specifically targets object hallucinations, and 2) it employs an object detector (Liu et al., 2023b) for hallucination detection (see Table 1).
Table 5 shows that the decoding-based methods are ineffective for detailed image captioning. Ironically, applying VCD significantly reduces the factuality of the LLaVA model while increasing coverage. Volcano yields only slight improvements in LLaVA’s captions. However, CapMAS substantially enhances the factuality of the captioning model compared to the other methods. These results suggest that methods proposed to enhance MLLM factuality should be evaluated not only on tasks requiring short responses, such as VQA, but also on detailed image captioning tasks.
@@ Appendix E Case
Appendix F Prompt Templates

### v3

@@ Abstract
Multimodal large language models (MLLMs) excel at generating highly detailed captions but often produce hallucinations. Our analysis reveals that existing hallucination detection methods struggle with detailed captions. We attribute this to the increasing reliance of MLLMs on their generated text, rather than the input image, as the sequence length grows. To address this issue, we propose a multiagent approach that leverages LLM-MLLM collaboration to correct given captions. Additionally, we introduce an evaluation framework and a benchmark dataset to facilitate the systematic analysis of detailed captions. Our experiments demonstrate that our proposed evaluation method better aligns with human judgments of factuality than existing metrics and that existing approaches to improve the MLLM factuality may fall short in hyper-detailed image captioning tasks. In contrast, our proposed method significantly enhances the factual accuracy of captions, even improving those generated by GPT-4V. Finally, we highlight a limitation of VQA-centric benchmarking by demonstrating that an MLLM’s performance on VQA benchmarks may not correlate with its ability to generate detailed image captions. Our code and data are available at https://github.com/adobe-research/CapMAS.
@@ 3.3 Evaluation Methods
To validate this evaluation method, we use the DOCCI dataset (Onoe et al., 2024), which contains human-annotated detailed image captions. Specifically, for each image in a subset of the dataset, we prepare the following four types of captions (details provided in Appendix G):
@@ 4.1 Experimental Setup
We employ our proposed factuality and coverage metrics, along with CLAIR, all based on GPT-4o, to evaluate the generated captions. To ensure robust evaluation, we summarize the captions (Ge et al., 2024) generated from five different prompts using LLaMA-3-8B. The only CapMAS hyperparameter, , is tuned on a validation set of five examples sampled from the DCI dataset (Urbanek et al., 2024). The prompt templates are provided in Appendix G.
@@ 4.4 Comparison with Other Methods
Various methods have been proposed to mitigate hallucinations in MLLMs, and they have primarily been validated on VQA and simple captioning benchmarks. We compare CapMAS with three recent decoding-based methods (VCD, OPERA, and SPARC (Jung et al., 2025)), two corrector-based methods (LURE and Volcano), and one training-based method (LRV) from the perspective of detailed image captioning. All methods, except for LRV and LURE, use LLaVA-v1.5-7B, while the LRV and LURE methods employ the MiniGPT-4 model (Zhu et al., 2023) as provided by their respective authors. For reference, VisualFactChecker (VFC) (Ge et al., 2024) is also a pipeline composed of pre-trained models that revise initial captions, similar to our approach. However, the inability to reproduce VFC, as its authors have not provided the necessary resources for reproduction, prevents a direct comparison with our method. Nonetheless, we can infer that our method outperforms VFC because 1) VFC specifically targets object hallucinations, and 2) it employs an object detector (Liu et al., 2023b) for hallucination detection (see Table 1).
Table 5 shows that the decoding-based methods are ineffective for detailed image captioning. Ironically, applying VCD significantly reduces the factuality of the LLaVA model while increasing coverage. SPARC does not show a notable impact in terms of factuality, but it proves to be considerably beneficial with respect to coverage. Volcano yields only slight improvements in LLaVA’s captions. However, CapMAS substantially enhances the factuality of the captioning model compared to the other methods. These results suggest that methods proposed to enhance MLLM factuality should be evaluated not only on tasks requiring short responses, such as VQA, but also on detailed image captioning tasks.
@@ 4.5 Consistency Between MLLM Captioning and VQA Evaluation results
We provide additional results on fluency, cost, and test-time scaling in Appendix F.
@@ 5 Conclusion
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2021-II211343, RS-2022-II220959, Artificial Intelligence Graduate School Program (Seoul National University)], the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720), a grant from Yang Young Foundation, and the BK21 FOUR program of the Education and the Research Program for Future ICT Pioneers, Seoul National University in 2025.
@@ Appendix E Case
Appendix F Additional Experiments
F.1 Fluency Analysis
We investigate the effect of our framework on the fluency of the generated captions. We employ an LLM-based evaluation to measure the basic quality of language in the captions generated by CapMAS. Specifically, we utilize GPT-4o with the following prompt:
You are a language expert evaluating the fluency of image captions.
Fluency refers to how grammatically correct, natural, and well-formed the text sounds to a native English speaker. A fluent caption should be grammatically correct, free of awkward phrasing, and read smoothly.
Evaluate the fluency of the following caption and return your output **strictly in JSON format** with:
- “reason”: a key reason for your scoring
- “score”: a number between 0 (completely disfluent) and 100 (perfect fluency)Caption: “{caption}”
In addition to evaluating the captions generated by CapMAS, we also assess human-written, detailed captions for the same set of images.
| Captions generated by | Fluency |
| Human | 89.0 |
| CapMAS (LLaVA-v1.5-7B) | 93.4 |
| CapMAS (LLaVA-NeXT-7B) | 93.4 |
| CapMAS (LLaVA-NeXT-13B) | 93.6 |
| CapMAS (InternVL-Chat-V1.5) | 94.1 |
The results in Table 10 demonstrate that the captions generated by CapMAS achieve even higher fluency scores than the human-generated captions. This can be attributed to the final stage of CapMAS, in which the corrector LLM helps preserve or even improve the fluency of the captions.
F.2 Effect of Self-Refine on Image Captioning
CapMAS can be interpreted as an inference-time scaling strategy. However, allocating more compute at inference time does not necessarily lead to better results. To demonstrate this, we test the effectiveness of inference-time scaling through Self-Refine (SR) (Madaan et al., 2023). Since validating and revising one’s output requires advanced reasoning capabilities, we conducted the experiments using GPT-4V. For SR, we used the following prompt:
You are given an image and its corresponding caption. Your task is to:
1. Analyze the image and compare it with the caption.
2. Identify and correct any factual or descriptive errors in the caption based on the image.
3. Refine the caption for clarity, correctness, and completeness — even if the original caption is mostly accurate.Show your reasoning and then provide a final refined caption.
| Method | CLAIR | Factuality | Coverage |
| Base | 82.4 | 77.1 | 53.5 |
| +SR (x1) | 79.9 | 72.3 | 50.4 |
| +SR (x2) | 79.1 | 70.6 | 50.1 |
| +SR (x3) | 78.5 | 69.8 | 49.7 |
Table 11 shows that having the model revise its own captions iteratively does not lead to better results.
F.3 Efficiency of CapMAS
CapMAS is an approach that achieves better results by incurring additional cost at inference time. Although it involves a multi-model pipeline, CapMAS can offer a better cost-performance trade-off than Self-Refine (SR) methods for the following reasons:
Most of CapMAS’s cost lies in the final step, where the corrector LLM generates a refined caption based on the initial caption and the True/False classification results of the propositions. The decomposition step involves shorter sequences, and the MLLM used for proposition classification can process them in parallel, generating only a single token (True or False) per proposition. SR processes long sequences that include the original caption, detailed feedback, and the refined caption. Considering the length and complexity of the feedback (Madaan et al., 2023), CapMAS and SR can be seen as comparable in cost.
CapMAS’s performance does not heavily depend on the capability of the LLM used. This suggests that the LLM-related cost in the pipeline could potentially be further reduced.
Appendix G Prompt Templates